Hi there, i am trying to find my average for a module at university.

i have two units for the module:
one worth 60% where i got a C12
and the other worth 40% where i got an A18

the marking scale is as follows:
D (9-11) C (12-14) B (15-17) A (18-20)


Answer by Theo(13342)

if i understand this correctly, you got a 12 and an 18.
since the 12 is worth 60% and the 18 is worth 40%, then your weighted score would be 60% * 12 + 40% * 18 = 14.4 which would put you in the C14 category if they categorize you to the nearest integer.
it appears they are scoring from 0 to 20 where 0 = 0 and 20 = 100%
this means that the ratio of score units to grade percentage points is equal to 5.
this means that if you score a 12, then your grade percentage for the test was 12 * 5 = 60, and if you score an 18, then your grade percentage for the test was 18 * 5 = 90.
60% * 60 and 40% * 90 equals an overall grade percentage of 72.
based on the ratio, an overall grade percentage of 72 results in a score of 14.4 units because 72/5 = 14.4.
that's what i think anyway.
your overall score would be 14.4 which puts you closer to C14 than to B15 which would round you into the C14 category.
that's what i think anyway.